A line contains the point (-3,-1) and has a slope of 1/3. Which equation represents this line?
A) x + 3y = 0
B) x - 3y = 0
C) 3x + y = 0
D) 3x - y = 0

Answer by lwsshak3(11628)

A line contains the point (-3,-1) and has a slope of 1/3. Which equation represents this line?
A) x + 3y = 0
B) x - 3y = 0
C) 3x + y = 0
D) 3x - y = 0
***
x-3y=0
3y=x
y=x/3
ans: B)